shift summary for owl shift 06/14/99

Sandrine Laplace (laplace@hep.uchicago.edu)
Mon, 14 Jun 1999 07:17:53 -0500

Shift Summary for owl shift, 06/14/99

On Shift: Sandrine,Hogan

also Elizabeth, Paty, Rick K, Sidney, Ed shows up at 6:45, just at time
for the pipeline problem.

Hours of beam delivered at all intensities: 6 hours
Hours of beam delivered at nominal intensity: 5 hours at 2e12
Hours of E832 data (special runs not included): 4 hours
Hours of special runs (laser scans, muon runs...): 1 hour

Accounting of lost time:

========================================================================
Runs ended this shift:
13583 - MU_4E12 - aborted due to overfull disk
13584 - BEAM_832 - Swapped magnet polarity to positive
13585 - MU_4E12 - Data for REG calibration
13586 - BEAM_832 - Didn't initialize properly: aborted.
13587 - BEAM_832 - still running at the end of the run.

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift: 1 access to check crate 6 of pipeline.

========================================================================
Overview of shift:

-- do muon runs to calibrate the regenerator.
-- Swap the magnet polarity (NM4AN) to positive, and then take data.

========================================================================
Description of Shift:

00:00 Inherit from previous shift with muons.
00:20 Starts run number 13583 with configuration BEAM_MU_4E12 in order
to do the regenerator calibration after a few changes. Data
written only on disk.
We are running 4e11/40 sec and 1 batch.
00:40 Run aborts because the disk ktev1a is overfull. By trying to save
the
histograms, we realize it doesn't work for a muon run. Rick K
fixes it.
1:00 We had trouble mounting QKB004. It doesn't have an End Of File
marker
according to Sydney, which could be due to a fresh_start_sequence
done in a middle of a run. We decide to overwrite the tapes since no
precious data have been taken (beam tests...).
1:15 Change NM4AN polarity to positive (ran AN832NEGTOP and AN832UP).
1:50 Start hadrons run number 13584 (BEAM_832) at 8.75e11, 3 batches.
1:55 Problem with the tape driver #5 of ktev3 : the cleaning light
still coming on even after we cleaned it. We mounted a new tape,
and it seems to work. The old tape was probably bad.
2:10 The beam intensity is up to 1.62e12.
2:20 Rick gives us a course about the BOA philosophical meaning, with
equations and drawings. Everybody is listenning in silence...
3:10 BOA is 2.06 (and now, I know what it means), Number of batches is
now 10.
3:30 Switch back to muons for the REG calibration.
3:42 Start muons run number 13585, run it for 10 minutes, which
is enough for REG calibration.
3:50 Switch back to hadrons.
4:00 Start BEAM_832 run number 13586.
Rick K compared acc. energy for run 13584 and run 9225 (97).
It turns out that the average accidental energy is 3 Gev higher in
99, because of a mysterious pic at ~3 GeV in the energy
distribution.
Sandrine is running a job to get the 99 accidental energy
illumination.
4:15 Run 13586 didn't initiate properly (kumquat problems), do a
fresh_start_sequence. Start run 13587,
5:05 Problems with slowdaq : restart process mon_caen.tcl on atev3
Then lost the spill number: restart all processes, which fix it.
5:20 I really need a coffee...
6:20 Problems with pipeline. Crate 6 don't want to be rebooted.
7:00 Ed, Hogan, Sandrine make access to check crate 6 down stairs.
It seem to be a power supplie problem. The up-stairs alarm for that
doesn't work. The problem remains at the end of the shift.

========================================================================
Items to Follow up on Next Shift

-- crate 6 pipeline problem.

Regards,

Sandrine and Hogan